### v1

@@ Abstract
Comet C/2025 K1 (ATLAS) reached perihelion at 0.33 au on 2025 October 8. Daily monitoring by the LCO Outbursting Objects Key Project revealed a major activity increase between November 2 and 4, accompanied by rapid changes in coma morphology. Serendipitous HST/STIS acquisition images obtained on November 8–10 captured the comet only days after this event and resolved five fragments, providing an early high-resolution view of a nucleus in the process of disruption. Fragment motions and morphologies indicate a hierarchical fragmentation sequence, including a slow secondary split of fragment B. Back-extrapolation shows that both the primary and secondary breakups preceded their associated photometric outbursts by roughly one to three days. This consistent lag, together with the appearance of thin, short-lived arclets around fragment A in the first HST epoch, suggests that freshly exposed interior material warms rapidly but requires time before dust can be released efficiently. These combined ground- and space-based observations provide rare, time-resolved constraints on the thermal and structural evolution of a fragmented comet near perihelion and highlight the scientific value of capturing a nucleus within days of disruption.
@@
Fragment kinematics and a consistent 1–3 day delay between breakup and brightening show that post-fragmentation activation is controlled by thermal diffusion into freshly exposed interior layers, analogous to delayed activity following Rosetta’s cliff collapses on 67P.
Short-lived, shell-like arclets observed around fragment A in the first HST epoch indicate dust-shell release during initial warm-up rather than gas-driven pressure waves, revealing the physical processes governing early post-fragmentation morphology.
@@ 1 Introduction
Following perihelion, C/2025 K1 was monitored closely by the LCO Outbursting Objects Key (LOOK) Project (Lister et al., 2022). Comparison of the observations on November 4 with ZTF photometry showed that the comet underwent at least a 0.9 mag outburst between November 2 and 4, accompanied by the rapid development of a two-component coma: an extended parabolic outer feature and a compact oblate inner coma of freshly released material (Kelley et al., 2025). Independent small-telescope photometry (A. Pearce, priv. comm.) places the onset between November 3 10:33 UT and November 4 10:33 UT, and over the following nights the central condensation evolved to show multiple brightness peaks as fragments drifted far enough apart to be spatially separated from the ground (Minor Planet Center, 2025a, b).
@@ 2 Observations
Figure 1: LOOK r-band photometry of C/2025 K1 (ATLAS) measured in a circular aperture of 10,000 km in radius (orange squares).HST/STIS photometry within a similarly-sized aperture were bootstrapped to the Nov. 10 LOOK data and are depicted by blue triangles. The inset panel presents the absolute magnitudes of the individual fragments obtained from HST/STIS (cf. Fig. 3): open triangles denote fragment I; black, open, and grey squares represent fragments II, IIa, and IIb, respectively; grey circles indicate fragment IV (magnitudes of IV are decreased by 2-2 mag to aid visualization).
2.1 LOOK
@@ 2.2 HST/STIS
The images used here were context frames acquired for a spectroscopic survey using STIS and COS on board HST (Program 18135, PI D. Bodewits). Each consists of a 20.5 s exposure obtained with the MIRVIS detector and the F28X50LP long-pass filter, whose throughput spans 5500–10,000 Å with an FWHM of 2693 Å [^1] https://hst-docs.stsci.edu/stisihb/chapter-14-imaging-reference-material/14-3-ccd/ccd-long-pass-imaging-f28x50lp. The point-spread function has $R_{80}\approx$ 0.15 arcsec. Three pairs of STIS acquisition images were obtained over a two-day interval shortly after the comet’s major outburst, providing three well-spaced epochs for tracking the motion, morphology, and evolving activity of the fragments. The observations span 31–33 days past perihelion and cover modest changes in heliocentric and geocentric distance as well as phase angle. A complete observing log, including times, geometry, and exposure details for all three epochs, is given in Table 1. These three epochs provide the basis for tracking the motion, morphology, and evolving activity of the fragments.
We measured the brightness of each fragment using a circular aperture with the STIS $R_{80}$ radius of 0.15′′. Photon count rates were converted to fluxes and magnitudes using the calibration constants in the FITS headers (Proffitt, 2006), without background subtraction. At the $\sim$0.6 au geocentric distance of the observations, the 0.15′′ PSF corresponds to a physical scale of $\sim$68 km—far larger than any plausible nucleus—so the extracted fluxes are dust dominated. We therefore applied the Schleicher–Marcus composite dust phase correction [^2] https://asteroid.lowell.edu/comet/dustphase/table to account for the large scattering angles. We also measured the total flux in each frame using a fixed 10,000 km-radius aperture centered on the image, again without background subtraction and with the same phase correction. A solar spectrum through the F28X50LP filter is comparable to an $r$-band magnitude, so no color correction was applied. To compare the STIS results with the nearly contemporaneous LOOK photometry on Nov. 10 (within 2.5 hr), we converted both to $H(1,0,0)$. The LOOK measurement is 0.5 mag brighter than the STIS result, which we attribute to differences in filter bandpasses. We therefore bootstrap the STIS photometry to the LOOK scale by subtracting 0.5 mag from all STIS magnitudes.
@@ 3 Results
Figure 3: HST/STIS MIRVIS images of C/2025 K1 (ATLAS) obtained on 2025 Nov. 8.56 (left), Nov. 9.82 (center), and Nov. 10.54 (right). Each panel uses a logarithmic stretch with the same physical and intensity scale. The field of view is $20\mbox{${}^{\prime\prime}$}\times 15\mbox{${}^{\prime\prime}$}$, corresponding to roughly $9{,}000\times 6{,}500$ km at the comet. Fragments are labeled following the notation introduced in the text: the two brightest components (I and II) are present in the first epoch, with fainter fragments III and IV located progressively farther along the projected orbit. By the second epoch, fragment II has separated into components IIa and IIb, and the relative motions and evolving morphologies of all fragments are evident, including arclets around I, a parabolic inner coma around II, the forward fan of IV, and the fading of III. In the final panel, the brightening and coma development of fragment IIb are clearly visible. Grayscale values represent counts in a 21.5 s exposure.
@@ 3.1 Ground-Based Photometry and Context
LOOK photometry for Nov. 4-–22, spanning the HST observing window, is shown in Fig. 1. The –0.9 mag outburst of (Kelley et al., 2025) is apparent as a decrease in the brightness with time starting with our first observation on November 4 12:16 UTC. Two additional outbursts are found in the data, on November 14 4:17 UTC and November 22 9:04 UTC, with strengths of $0.94\pm 0.08$ and $0.33\pm 0.06$ mag for our 10,000 km aperture. Figure 2 shows selected images of the comet around the time of the HST observations.
STIS photometry extracted within a 10,000 km aperture yield nearly constant absolute magnitudes of H=11.0H=11.011.111.1 across the three epochs (Nov. 8.56–10.54; Fig. 1). Although the STIS images show that fragment IIb was already brightening by Nov. 10.53, this increase was not immediately apparent in the LOOK photometry because IIa and IIb remained unresolved from the ground. Once their separation increased and IIb’s intrinsic activity strengthened, the effect became visible: between Nov. 12 12:01 (T-Tp = 35.06 days) and Nov. 14 05:23 (T-Tp = 36.78 days), the LOOK data suggest that fragment IV underwent an $\sim$1.0 mag outburst, producing a newly pronounced parabolic coma clearly distinguishable from the pre-existing coma. This represents the integrated ground-based signature of activity that began at least two days earlier, as captured directly by HST.
A later episode of activity is visible in the LOOK monitoring near the end of the observing window. A new outburst of the eastern fragment occurred between Nov. 21 09:57 and Nov. 22 09:04 UTC, as seen in consecutive r-band frames (logarithmic stretch) obtained on Nov. 20, 21, and 22. The Nov. 22 image shows a newly developed, compact parabolic coma coincident with the eastern fragment, distinct from the broader, fading dust structures associated with the earlier activity. By this date, fragment I had become increasingly diffuse, consistent with its eventual dispersal into a fading cloud rather than a stable condensation.
@@ 3.2 High-Resolution Morphology and Evolution of the Fragments
HST observations directly resolve the evolving activity into individual fragments (Fig. 3). The first STIS image, acquired on 2025 Nov. 8.56, reveals at least four fragments: two brighter components (I and II) separated by 3.4′′, and two fainter components (III and IV) located 4.4′′ and 5.1′′ ahead of I along the projected orbit. Fragment IV shows a distinct forward fan with an opening angle of roughly $100^{\circ}$, while fragment I exhibits arclets nearly perpendicular to the dust-tail direction. Fragment II appears to be the most active, surrounded by a compact parabolic coma.
In the second epoch (Nov. 9.82; 30.33 hr later), the I–II separation increased to 3.9′′, and II split into two components, IIa and IIb, separated by 0.24′′. Fragment III evolved from a compact clump into a diffuse smear, while IV remained clearly detectable and had moved to 6.2′′ ahead of I.
By the final HST epoch (Nov. 10.53; 17.22 hr later), the separation IIa-IIb increased to 0.30′′, the separation I–IIa had increased to 4.2′′, III had faded below detection, and IV appeared slightly diminished but still present at 6.7′′. Notably, IIb had brightened substantially, indicating the onset of significantly enhanced activity late in the sequence.
The arclets surrounding fragment I provide an additional diagnostic of its recent activity. Arclet-like features have been reported in other comets (e.g., Harris et al., 1997; Boehnhardt, 2004), but in C/2025 K1 they are seen only in the first HST epoch (Nov. 8.56). Within that epoch they remain tightly centered on A, show no measurable outward expansion over the paired frames, and appear as thin, shell-like structures nearly perpendicular to the dust-tail direction. This behavior is inconsistent with the global pressure-wave model proposed by Farnham et al. (2001) for C/1996 B2 (Hyakutake), and instead strongly suggests confined dust shells released when freshly exposed material was first illuminated and warmed after the primary breakup.
@@ 3.3 Fragment Sizes, Separation Velocities, and the Fragmentation Sequence
Using 0.15′′ (68 km) apertures, the scattering cross sections of the fragments show clear and systematic evolution across the three HST epochs. On Nov. 8.56, fragments I and II dominate with cross sections of 3.6 km2 and 9.4 km2, respectively, while fragment IV is much smaller, contributing only 0.5 km2. By the second epoch (Nov. 9.82), fragment II has divided into two pieces: IIa and IIb exhibit comparable cross sections of 6.0 km2 and 5.3 km2. Fragment I temporarily brightens to 7.2 km2, consistent with the development of arclets, while fragment IV remains faint at 0.6 km2. By the final epoch (Nov. 10.54), I settles to 5.6 km2, IIa remains steady at 5.6 km2, and IIb brightens markedly to 8.1 km2, mirroring the onset of enhanced activity visible in the morphology. Fragment IV is nearly unchanged at 0.6 km2. Interpreting these cross sections as upper limits to the effective radii (due to coma contamination) yields roughly 1.1 km for I, $\sim$1.7 km for II, $\sim$1.3 km for IIa and IIb, and $\sim$0.4 km for fragment IV.
@@
Extrapolating the projected I–II separation backwards places the primary fragmentation event near 2025 Nov. 1 (within several days), consistent with the timing of the 0.9 mag outburst observed between Nov. 2 and 4. Fragments III and IV must have separated earlier and at higher speeds, consistent with their smaller sizes and rapid fading. The secondary breakup of II into IIa and IIb is tightly constrained to the interval between Nov. 8.56 and 9.82, preceding the sharp brightening of IIb that becomes prominent in the final HST frame and in the LOOK survey several days later.
The timing of the morphological changes and the lightcurve implies a 1–3 day delay between fragment exposure and peak dust production. The shell-like arclets around fragment I reinforce this scenario: they are confined, stationary dust shells produced when freshly exposed surfaces first warmed and began releasing material, consistent with a thermal timescale of several rotation cycles.
We briefly explored an alternative description in which the separations are fit by constant accelerations rather than constant projected speeds. This produces a self-consistent single fragmentation epoch around 2025 Oct. 20, with higher accelerations for the faint fragments III and IV and lower accelerations for II. This date is consistent, within uncertainties, with an earlier outburst reported by Qicheng Zhang [^3] https://cometary.org/@qicheng/statuses/01K8BMPX63NCB3K0NXZMYF2XQW before post-perihelion LOOK monitoring resumed. However, the limited number of HST epochs, the reliance on projected distances rather than true 3D positions, and the observed secondary breakup of II argue that hierarchical fragmentation provides a more natural and physically plausible description. For these reasons, we adopt the simpler interpretation based on projected separation rates.

### v2

@@ Abstract
Comet C/2025 K1 (ATLAS) reached perihelion at 0.33 au on 2025 October 8. Daily monitoring by the LCO Outbursting Objects Key Project revealed a major activity increase between November 2 and 4, accompanied by rapid changes in coma morphology. Serendipitous HST/STIS acquisition images obtained on November 8–10 captured the comet only days after this event and resolved five fragments, providing an early high-resolution view of a nucleus in the process of disruption. Fragment motions and morphologies indicate a hierarchical fragmentation sequence, including a slow secondary split of fragment II. Back-extrapolation shows that both the primary and secondary breakups preceded their associated photometric outbursts by roughly one to three days. This consistent lag, together with the appearance of thin, short-lived arclets around fragment I in the first HST epoch, suggests that freshly exposed interior material warms rapidly but requires time before dust can be released efficiently. Given the comet’s close perihelion passage, rotational instability driven by enhanced outgassing torques is a plausible contributor to nucleus disintegration and dust release, and may represent the primary source of the observed brightening. These combined ground- and space-based observations provide rare, time-resolved constraints on the thermal and structural evolution of a fragmented comet near perihelion and highlight the scientific value of capturing a nucleus within days of disruption when thermal adjustment, dust mantle re-formation, and outgassing-driven torques jointly govern the onset of activity.
@@
Fragment kinematics and a consistent 1–3 day delay between breakup and brightening show that post-fragmentation activation is controlled by thermal adjustment of newly exposed surfaces, analogous to delayed activity following Rosetta’s cliff collapses on 67P.
Short-lived, shell-like arclets observed around fragment I in the first HST epoch indicate dust-shell release during initial warm-up rather than gas-driven pressure waves, revealing the physical processes governing early post-fragmentation morphology and the onset of efficient dust production.
@@ 1 Introduction
Following perihelion, C/2025 K1 was monitored closely by the Las Cumbres Observatory Outbursting Objects Key (LOOK) Project (Lister et al., 2022). Comparison of the observations on November 4 with Zwicky Transient Facility (ZTF) photometry showed that the comet underwent at least a 0.9 mag outburst between November 2 and 4, accompanied by the rapid development of a two-component coma: an extended parabolic outer feature and a compact oblate inner coma of freshly released material (Kelley et al., 2025). Independent small-telescope photometry (A. Pearce, priv. comm.) places the onset between November 3 10:33 UT and November 4 10:33 UT, and over the following nights the central condensation evolved to show multiple brightness peaks as fragments drifted far enough apart to be spatially separated from the ground (Minor Planet Center, 2025a, b).
@@ 2 Observations
Figure 1: LOOK r-band photometry of C/2025 K1 (ATLAS) measured in a circular aperture of 10,000 km in radius (orange squares).HST/STIS photometry within a similarly-sized aperture were bootstrapped to the Nov. 10 LOOK data and are depicted by blue triangles. The inset panel presents the absolute magnitudes of the individual fragments a within apertures with radii of 0.15 arcsec obtained from HST/STIS (cf. Fig. 3): open triangles denote fragment I; black, open, and grey squares represent fragments II, IIa, and IIb, respectively; grey circles indicate fragment IV (magnitudes of IV are decreased by 2-2 mag to aid visualization) The time on the x-axis is the time since perihelion, which occurred on Oct 8.4, 2025.
2.1 LCO Outbursting Objects Key Project
@@ 2.2 HST/STIS
The images used here were context frames acquired for a spectroscopic survey using the Space Telescope Imaging Spectrograph (STIS) and the Cosmic Origins Spectrograph (COS) on board HST (Program 18135, PI D. Bodewits). Each consists of a 20.5 s exposure obtained with the MIRVIS detector and the F28X50LP long-pass filter, whose throughput spans 5500–10,000 Å with an FWHM of 2693 Å [^1] https://hst-docs.stsci.edu/stisihb/chapter-14-imaging-reference-material/14-3-ccd/ccd-long-pass-imaging-f28x50lp. The radius of the point-spread function that encloses 80% of the flux ($R_{80})$$is\approx$ 0.15 arcsec. Three pairs of MIRVIS images were obtained over a two-day interval shortly after the comet’s major outburst, providing three well-spaced epochs for tracking the motion, morphology, and evolving activity of the fragments. The observations span 31–33 days past perihelion and cover modest changes in heliocentric and geocentric distance as well as phase angle. A complete observing log, including times, geometry, and exposure details for all three epochs, is given in Table 1. These three epochs provide the basis for tracking the motion, morphology, and evolving activity of the fragments.
We measured the brightness of each fragment using a circular aperture with the STIS $R_{80}$ radius of 0.15′′. Photon count rates were converted to fluxes and magnitudes using the calibration constants in the FITS headers (Proffitt, 2006), without background subtraction. At the $\sim$0.6 au geocentric distance of the observations, the 0.15′′ PSF corresponds to a physical scale of $\sim$68 km—far larger than any plausible nucleus—so the extracted fluxes are dust dominated. We therefore applied the Schleicher–Marcus composite dust phase correction [^2] https://asteroid.lowell.edu/comet/dustphase/table to account for the large scattering angles. We also measured the total flux in each frame using a fixed 10,000 km-radius aperture centered on the image, again without background subtraction and with the same phase correction. A solar spectrum through the F28X50LP filter is comparable to an $r$-band magnitude, so no color correction was applied. To compare the STIS results with the nearly contemporaneous LOOK photometry on Nov. 10 (within 2.5 hr), we converted both to absolute magnitudes $H(1,0,0)$. The LOOK measurement is 0.5 mag brighter than the STIS result, which we attribute to differences in filter bandpasses. We therefore bootstrap the STIS photometry to the LOOK scale by subtracting 0.5 mag from all STIS magnitudes.
@@ 3 Results
Figure 3: HST/STIS MIRVIS images of C/2025 K1 (ATLAS) obtained on 2025 Nov. 8.56 (T-Tp = 32.16 days) (left), Nov. 9.82 (T-Tp = 33.42 days) (center), and Nov. 10.54 (T-Tp = 34.14 days) (right). Each panel uses a logarithmic stretch with the same physical and intensity scale. The field of view is $20\mbox{${}^{\prime\prime}$}\times 15\mbox{${}^{\prime\prime}$}$, corresponding to roughly $9{,}000\times 6{,}500$ km at the comet. Fragments are labeled following the notation introduced in the text: the two brightest components (I and II) are present in the first epoch, with fainter fragments III and IV located progressively farther along the projected orbit. By the second epoch, fragment II has separated into components IIa and IIb, and the relative motions and evolving morphologies of all fragments are evident, including arclets around I, a parabolic inner coma around II, the forward fan of IV, and the fading of III. In the final panel, the brightening and coma development of fragment IIb are clearly visible. Grayscale values represent counts in a 21.5 s exposure.
@@ 3.1 Ground-Based Photometry and Context
LOOK photometry for Nov. 4–22 [T-Tp = 27.61–45.98 days], spanning the HST observing window, is shown in Fig. 1. The –0.9 mag outburst of (Kelley et al., 2025) is apparent as a decrease in the magnitude with time starting with our first observation on November 4 12:16 UTC[T-Tp = 27.61 days]. Two additional outbursts are found in the data, on November 14 04:17 UTC [T-Tp = 36.78 days] and November 22 09:04 UTC [T-Tp = 45.98 days], with strengths of $0.94\pm 0.08$ and $0.33\pm 0.06$ mag for our 10,000 km aperture. Figure 2 shows selected images of the comet around the time of the HST observations. After the outburst, the comet faded from $H(1,1,0)=10.9$ to 11.0 by November 10 [T-Tp = 33.56 days], with small deviations attributable to fragment brightening resolved by HST.
STIS photometry extracted within a 10,000 km aperture yield nearly constant absolute magnitudes of H=11.0H=11.011.111.1 across the three epochs (Nov. 8.56–10.54 [T-Tp = 32.16–34.14 days]; Fig. 1). Although the STIS images show that fragment IIb was already brightening by Nov. 10.53 [T-Tp = 34.13 days], this increase was not immediately apparent in the LOOK photometry because IIa and IIb remained unresolved from the ground. Once their separation increased and IIb’s intrinsic activity strengthened, the effect became visible: between Nov. 12 12:01 [T-Tp = 35.06 days] and Nov. 14 05:23 [T-Tp = 36.78 days], the LOOK data suggest that fragment IV underwent an $\sim$1.0 mag outburst, producing a newly pronounced parabolic coma clearly distinguishable from the pre-existing coma. This represents the integrated ground-based signature of activity that began at least two days earlier, as captured directly by HST.
A later episode of activity is visible in the LOOK monitoring near the end of the observing window. A new outburst of the eastern fragment occurred between Nov. 21 09:57 [T-Tp = 45.03 days] and Nov. 22 09:04 UTC [T-Tp = 45.98 days], as seen in consecutive r-band frames (logarithmic stretch) obtained on Nov. 20, 21, and 22 [T-Tp = 44.03–46.03 days]. The Nov. 22 image shows a newly developed, compact parabolic coma coincident with the eastern fragment, distinct from the broader, fading dust structures associated with the earlier activity. By this date, fragment I had become increasingly diffuse, consistent with its eventual dispersal into a fading cloud rather than a stable condensation.
@@ 3.2 High-Resolution Morphology and Evolution of the Fragments
HST observations directly resolve the evolving activity into individual fragments (Fig. 3). The first STIS image, acquired on 2025 Nov. 8.56 [T-Tp = 32.16 days], reveals at least four fragments: two brighter components (I and II) separated by 3.4′′, and two fainter components (III and IV) located 4.4′′ and 5.1′′ ahead of I along the projected orbit. Fragment IV shows a distinct forward fan with an opening angle of roughly $100^{\circ}$, while fragment I exhibits arclets nearly perpendicular to the dust-tail direction. Fragment II appears to be the most active, surrounded by a compact parabolic coma.
In the second epoch (Nov. 9.82 [T-Tp = 33.42 days]; 30.33 hr later), the I–II separation increased to 3.9′′, and II split into two components, IIa and IIb, separated by 0.24′′. Fragment III evolved from a compact clump into a diffuse smear, while IV remained clearly detectable and had moved to 6.2′′ ahead of I. By the final HST epoch (Nov. 10.53 [T-Tp = 34.13 days]; 17.22 hr later), the separation IIa–IIb increased to 0.30′′, the separation I–IIa had increased to 4.2′′, III had faded below detection, and IV appeared slightly diminished but still present at 6.7′′. Notably, IIb had brightened substantially, indicating the onset of significantly enhanced activity late in the sequence.
The arclets surrounding fragment I provide an additional diagnostic of its recent activity. Arclet-like features have been reported in other comets (e.g., Harris et al., 1997; Boehnhardt, 2004), but in C/2025 K1 they are seen only in the first HST epoch (Nov. 8.56 [T-Tp = 32.16 days]). Within that epoch they remain tightly centered on I, show no measurable outward expansion over the paired frames, and appear as thin, shell-like structures nearly perpendicular to the dust-tail direction. This behavior is inconsistent with the global pressure-wave model proposed by Farnham et al. (2001) for C/1996 B2 (Hyakutake), and instead strongly suggests confined dust shells released when freshly exposed material was first illuminated and began to warm, or when a newly formed dust mantle underwent rapid thermal fracturing, an origin fully consistent with the thermal-delay scenario inferred from the lightcurve.
@@ 3.3 Fragment Sizes, Separation Velocities, and the Fragmentation Sequence
Using 0.15′′ (68 km) apertures, the scattering cross sections of the fragments show clear and systematic evolution across the three HST epochs. On Nov. 8.56 [T-Tp = 32.16 days], fragments I and II dominate with cross sections of 3.6 km2 and 9.4 km2, respectively, while fragment IV is much smaller, contributing only 0.5 km2. By the second epoch (Nov. 9.82 [T-Tp = 33.42 days]), fragment II has divided into two pieces: IIa and IIb exhibit comparable cross sections of 6.0 km2 and 5.3 km2. Fragment I temporarily brightens to 7.2 km2, consistent with the development of arclets, while fragment IV remains faint at 0.6 km2. By the final epoch (Nov. 10.54 [T-Tp = 34.14 days]), I settles to 5.6 km2, IIa remains steady at 5.6 km2, and IIb brightens markedly to 8.1 km2, mirroring the onset of enhanced activity visible in the morphology. Fragment IV is nearly unchanged at 0.6 km2. Interpreting these cross sections as upper limits to the effective radii using the assumptions listed in Graykowski and Jewitt (2019) yields roughly 1.1 km for I, $\sim$1.7 km for II, $\sim$1.3 km for IIa and IIb, and $\sim$0.4 km for fragment IV.
@@
Extrapolating the projected I–II separation backwards places the primary fragmentation event near 2025 Nov. 1 [T-Tp = 23.56 days], consistent with the timing of the 0.9 mag outburst observed between Nov. 2 and 4 [T-Tp = 24.56–27.61 days]. Fragments III and IV must have separated earlier and at higher speeds, consistent with their smaller sizes and rapid fading. The secondary breakup of II into IIa and IIb is tightly constrained to the interval between Nov. 8.56 and 9.82 [T-Tp = 32.16–33.42 days], preceding the sharp brightening of IIb that becomes prominent in the final HST frame and in the LOOK survey several days later.
The timing of the morphological changes and the lightcurve implies a 1–3 day delay between fragment exposure and peak dust production. The shell-like arclets around fragment I reinforce this scenario, consistent with a thermal timescale of several rotation cycles.
We briefly explored an alternative description in which the separations are fit by constant accelerations rather than constant projected speeds. This produces a self-consistent single fragmentation epoch around 2025 Oct. 20 [T-Tp = 11.60 days], with higher accelerations for the faint fragments III and IV and lower accelerations for II. This date is consistent, within uncertainties, with an earlier outburst reported by Qicheng Zhang [^3] https://cometary.org/@qicheng/statuses/01K8BMPX63NCB3K0NXZMYF2XQW before post-perihelion LOOK monitoring resumed. However, the limited number of HST epochs, the reliance on projected distances rather than true 3D positions, and the observed secondary breakup of II argue that hierarchical fragmentation provides a more natural and physically plausible description. For these reasons, we adopt the simpler interpretation based on projected separation rates.